OpenStudy (abbyselena2199):

HELP!!!!! The Smiths spent $2,000 on clothes in a particular year. If they increase the amount they spend on clothes by 10% each year, the Smiths will spend a total of $ on clothing in 4 years.

rebeccaxhawaii (rebeccaxhawaii):

WELCOME TO OS!!!

rebeccaxhawaii (rebeccaxhawaii):

do you have the fomula you would use?

OpenStudy (abbyselena2199):

I dont know lol i just really need help.

OpenStudy (welshfella):

can you work out what 10% of $2000 is?

OpenStudy (abbyselena2199):

its 200 right?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

so in the second year thwy have 2000 + 200 = $2200 to spend

OpenStudy (abbyselena2199):

Ok?

OpenStudy (welshfella):

and in the third year its 10% of 2200 Can you work that out?

OpenStudy (abbyselena2199):

I know what your gonna say after i get all the years ill need to add them together well the total would be 10,164 but my plato says its wrong

OpenStudy (abbyselena2199):

:(

OpenStudy (welshfella):

2000 2200 2200 * 1.1 = 2420 2420 * 1.1 = 2662 adding these up gives you $9282

OpenStudy (abbyselena2199):

I've tried that answer also. It also says it's wrong.

OpenStudy (welshfella):

9282 is correct

OpenStudy (abbyselena2199):

Okay. I'll try it again.

OpenStudy (welshfella):

I dont know if you've studied sequances but this is a geometric sequence with first term 2000 and common ratio 1.1 and sum for 4 years = 2000* ( 1.1^4 - 1) ---------- 1.1 - 1 which comes to 9282

OpenStudy (abbyselena2199):

Wanna help with another one?

OpenStudy (welshfella):

ok but first tell me what does xP mean in chat?

OpenStudy (welshfella):

(unless its something rude!!)